### v1

# Appendix 0.A Quantitative Results

Appendix
Appendix 0.A Quantitative Results
We compute several metrics for 3D generation results with SDS[poole2022dreamfusion, wang2023score] and FSD. The hyperparameters are kept the same as those in Appendix 0.B. When using Stable Diffusion[song2020denoising] as backbone, we use 16 prompts and 4 random seeds for each prompt (Tab. 1). When using MVDream[shi2023mvdream] as backbone, we also use 16 prompts and 4 random seeds for each prompt (Tab. 2).
Table 1: Stable Diffusion[song2020denoising] as backbone
Method CLIP ($\uparrow$) IS ($\uparrow$) CROSS-FID ($\uparrow$)
DDIM Images $33.72\pm 1.83$ $1.68\pm 0.55$
SDS[poole2022dreamfusion, wang2023score] $32.57\pm 1.43$ $1.58\pm 0.47$ $106.5\pm 58.3$
FSD(ours) $\mathbf{32.72\pm 1.56}$ $\mathbf{1.78\pm 0.49}$ $\mathbf{141.8\pm 57.9}$
Table 2: MVDream[shi2023mvdream] as backbone
Method CLIP ($\uparrow$) IS ($\uparrow$) CROSS-FID ($\uparrow$)
DDIM Images $34.64\pm 2.56$ $2.02\pm 0.47$
SDS[poole2022dreamfusion, wang2023score] $\mathbf{30.93\pm 3.40}$ $1.77\pm 0.37$ $86.6\pm 33.4$
FSD(ours) $30.12\pm 3.07$ $\mathbf{2.13\pm 0.35}$ $\mathbf{174.8\pm 44.5}$
0.A.0.1 CLIP score
We compute CLIP score[hessel2021clipscore, radford2021learning] using ViT-B/32 to measure the semantic similarity between the renderings of the generated 3D object and the input text prompt. We sample 24 views for each prompt and each seed when computing CLIP score.
0.A.0.2 IS score
We compute IS score[salimans2016improved] to measure both the image quality and diversity. We first compute the IS scores of sampled views for each prompt and then average the IS scores across prompts.
0.A.0.3 CROSS-FID score
To directly measure the diversity of generation results, it is natural to measure the inception distance between different generated samples. We first sample 24 views for each prompt and each seed. Then we separate the images corresponding to random seeds 0, 1 and 2, 3 into two sets of images. We compute FID[heusel2017gans] of the two sets of images and average the FID score across prompts. We term this score as CROSS-FID score since it is different from the standard way of using FID to evaluate GANs[heusel2017gans].
Appendix 0.B Additional Experiment Results
0.B.1 Additional Generation Results
Figure 1: Comparisons between FSD and SDS-like baseline methods. We present results of FSD and compare them to baseline methods, namely ProlificDreamer (VSD)[wang2024prolificdreamer], SDS[poole2022dreamfusion, wang2023score], CSD[ho2022classifier]. We use random seeds from 0 to 3 for each prompt.
Figure 2: Experiment results of FSD and SDS[poole2022dreamfusion, song2020score] with Stable Diffusion[song2020denoising] as backbone.
Figure 3: Experiment results of FSD and SDS[poole2022dreamfusion, song2020score] with MVDream[shi2023mvdream] as backbone.
We provide additional 3D experiment results using FSD and other SDS-like methods in Fig. 1, Fig. 2 and Fig. 3.
Specifically, in Fig. 1, we present results of FSD and compare them to baseline methods, namely ProlificDreamer (VSD)[wang2024prolificdreamer], SDS[poole2022dreamfusion, wang2023score], CSD[ho2022classifier]. In Fig. 2, we provide additional comparisons between SDS and FSD with Stable Diffusion[song2020denoising] as the backbone Diffusion Model. In Fig. 3, we provide additional comparisons between SDS and FSD with MVDream[shi2023mvdream] as the backbone Diffusion Model. We do not provide experiment results with Diffusion Models designed for novel view synthesis (e.g. SyncDreamer[liu2023syncdreamer]) since we found FSD does not improve the diversity of the back view of an object significantly given the front view when using Diffusion Models of such category.
0.B.2 Additional Ablations
$\Theta=120^{\circ}$
$\Theta=180^{\circ}$
$\Theta=240^{\circ}$
$\Theta=300^{\circ}$
$\Theta=360^{\circ}$
Figure 4: Ablation on other hyperparameters in $\boldsymbol{\epsilon}(\boldsymbol{c})$. We use a parameter $\Theta$ to control the size of noise world map ($H$ and $W$) in $\boldsymbol{\epsilon}(\boldsymbol{c})$. When $\Theta$ is larger, the “radius $r_{+}$” (Eq. 13 and Eq. 14) of the noise world map is smaller and FSD trends to generate smaller 3D models. In practice, we found FSD is prone to the parameter $\Theta$.
We use a parameter $\Theta$ to control the noise world-map’s size ($H$ and $W$). $H$ and $W$ are determined by $\Theta$ according to $H=H_{\text{hidden}}\frac{\pi}{\Theta}$ and $W=W_{\text{hidden}}\frac{2\pi}{\Theta}$. We visualize the results corresponding to different $\Theta$s in Fig. 4.
(a) vanilla constant noise function
(b) world-map noise function (proposed)
Figure 5: Ablation on design of $\boldsymbol{\epsilon}(\boldsymbol{c})$. We use noise function $\boldsymbol{\epsilon}(\boldsymbol{c})$ to align noise prior in FSD. We present RGB maps and depth maps of the generated samples using random seeds from 0 to 3. A vanilla constant function may result in holes on the surfaces. In contrast, the world-map noise function will not lead to any obvious flaws.
We also provide an additional comparison between the vanilla constant noise function and the world-map noise function in Fig. 5. We find the constant noise function may harm the geometry of the generation results.
0.B.3 Additional Experiments on Noise Prior
Figure 6: Impact of initial noise $\tilde{\boldsymbol{\epsilon}}$. We provide additional 2D experiment results generated by FSD show the impact of the initial noise $\tilde{\boldsymbol{\epsilon}}$. The patches framed with the same color in the same row share the same initial noise patches.
We provided additional generation results of FSD on 2D image space with shuffled or flipped initial noise in Fig. 6. The patches framed with the same color in the same row share the same initial noise patches $\tilde{\boldsymbol{\epsilon}}$.
Appendix 0.C Implementation Details
0.C.1 3D Experiments using FSD
The backbone Diffusion Model for Fig. 1 in the main text is MVDream[shi2023mvdream]. The configuration for this figure is the same as Fig. 7 in the main text.
We use sqrt-annealing proposed by HiFA[zhu2023hifa] for 3D experiments in this work. We use the same CFG[ho2022classifier] scale for SDS[shi2023mvdream, wang2023score] and FSD. We use CFG=1 when applying image loss trick[zhu2023hifa] since we find the reconstructed images are unrealistic when the CFG scale is large, which is not good guidance. We follow the default setting of threestudio[threestudio2023] code base for other hyperparameters. We keep the iteration number and learning rate of FSD the same as SDS[shi2023mvdream, wang2023score] in 3D experiments. We mainly conduct our experiments using NeRF representation[muller2022instant] since SDS-like methods are not sensitive to the form of 3D representations.
0.C.2 2D Experiments using FSD
Table 3: Implementation details on 2D experiments with FSD.
Methods Name SDS[poole2022dreamfusion, wang2023score] NFSD[katzir2023noise] VSD[wang2024prolificdreamer] FSD(ours) DDIM[song2020denoising]
Iteration Num 500 500 500 500 50
CFG[ho2022classifier] 100 7.5 7.5 7.5 7.5
Learning Rate 2e-2 2e-2 2e-2 3e-3 -
Optimizer Adam[kingma2014adam] Adam Adam Adam -
Timestep Annealing linear linear linear linear -
Here we describe the details of the experiments on 2D images with FSD in the main text. For Fig. 3 in the main text, we show the implementation details in Tab. 3. Below, we provide the loss functions for convenient reference.
Let denote the prediction of Diffusion Models as $\boldsymbol{\epsilon}_{y}^{t}=\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t}|y,t)$ and $c$ the classifier-free-guidance (CFG)[ho2022classifier] scale. Then
$\nabla_{\theta}L_{\text{SDS}}^{\theta}=\mathbb{E}_{\boldsymbol{\epsilon},\boldsymbol{c},t}\left[\left(c\cdot(\boldsymbol{\epsilon}^{t}_{y}-\boldsymbol{\epsilon}^{t}_{\emptyset})+(\boldsymbol{\epsilon}^{t}_{\emptyset}-\boldsymbol{\epsilon})\right)\frac{\partial\boldsymbol{g}_{\theta}(\boldsymbol{c})}{\partial\theta}\right]$ (1)
is the loss function of SDS[poole2022dreamfusion, wang2023score], where $\emptyset$ is the empty prompt.
$\nabla_{\theta}L_{\text{NFSD}}^{\theta}=\mathbb{E}_{\boldsymbol{\epsilon},\boldsymbol{c},t}\left[\left(c\cdot(\boldsymbol{\epsilon}^{t}_{y}-\boldsymbol{\epsilon}^{t}_{\emptyset})+(\boldsymbol{\epsilon}^{t}_{\emptyset}-\boldsymbol{\epsilon}^{t}_{y_{\text{neg}}})\right)\frac{\partial\boldsymbol{g}_{\theta}(\boldsymbol{c})}{\partial\theta}\right]$ (2)
is the loss function of NFSD[katzir2023noise], where $y_{\text{neg}}$ is a text negative prompt.
$\nabla_{\theta}L_{\text{VSD}}^{\theta}=\mathbb{E}_{\boldsymbol{\epsilon},\boldsymbol{c},t}\left[\left(c\cdot(\boldsymbol{\epsilon}^{t}_{y}-\boldsymbol{\epsilon}^{t}_{\emptyset})+(\boldsymbol{\epsilon}^{t}_{\emptyset}-\boldsymbol{\epsilon}^{t}_{\text{lora}})\right)\frac{\partial\boldsymbol{g}_{\theta}(\boldsymbol{c})}{\partial\theta}\right]$ (3)
is the loss function of VSD[wang2024prolificdreamer], where $\boldsymbol{\epsilon}^{t}_{\text{lora}}$ the Diffusion Model fine-tuned by LoRA[hu2021lora].
$\nabla_{\theta}L_{\text{FSD}}^{\theta}=\mathbb{E}_{\boldsymbol{c},t}\left[\left(c\cdot(\boldsymbol{\epsilon}^{t}_{y}-\boldsymbol{\epsilon}^{t}_{\emptyset})+(\boldsymbol{\epsilon}^{t}_{\emptyset}-\boldsymbol{\epsilon}(\boldsymbol{c}))\right)\frac{\partial\boldsymbol{g}_{\theta}(\boldsymbol{c})}{\partial\theta}\right]$ (4)
is the loss function of FSD. Additionally, $\boldsymbol{x}_{t}=\alpha_{t}\boldsymbol{g}_{\theta}(\boldsymbol{c})+\sigma_{t}\boldsymbol{\epsilon}(\boldsymbol{c})$ for FSD.
Our re-implementation of $L_{NFSD}$ is slightly different from the original[katzir2023noise] design of NFSD by ignoring the condition when $t<200$, to keep the formulation simple. In this way, we find $\boldsymbol{\epsilon}^{t}_{\text{lora}}$ used in VSD[wang2024prolificdreamer] may work in a similar way as $\boldsymbol{\epsilon}^{t}_{y_{\text{neg}}}$ in NFSD[katzir2023noise]. As a result, single-particle VSD may not be able to generate diverse samples since it is still mode-seeking, aligning with the observation of ESD[wang2023taming]. But we do not conduct further investigations which are out of the scope of our work.
Appendix 0.D Theory of Flow Score Distillation
We will provide some additional preliminaries and proof of Proposition 1 in the main text in this section.
0.D.1 Diffusion PF-ODE
The Diffusion PF-ODE [song2020score] can be written in the following form:
$\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{d}t}=f(t)\boldsymbol{x}_{t}-\frac{1}{2}g^{2}(t)\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x}_{t}|y),\quad\boldsymbol{x}_{T}\sim p_{T}(\boldsymbol{x}_{T}),$ (5)
where $f(t)=\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t},\;g^{2}(t)=\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma^{2}_{t}$, according to DPM-solver[lu2022dpm]. To get Eq. (5) in the main text, we take the derivative of $\boldsymbol{x}_{t}/\alpha_{t}$:
$\frac{\mathrm{d}(\boldsymbol{x}_{t}/\alpha_{t})}{\mathrm{d}t}=$ $\frac{1}{\alpha_{t}}\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{d}t}-\frac{\boldsymbol{x}_{t}}{\alpha_{t}^{2}}\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t}$ (6)
$=$ $\frac{1}{\alpha_{t}}\frac{\mathrm{d}\boldsymbol{x}_{t}}{\mathrm{d}t}-\frac{\boldsymbol{x}_{t}}{\alpha_{t}}f(t)$
$=$ $-\frac{g^{2}(t)}{2\alpha_{t}}\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x}_{t}|y)$
$=$ $-\frac{\frac{\mathrm{d}\sigma_{t}^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha_{t}}{\mathrm{d}t}\sigma^{2}_{t}}{2\alpha_{t}}\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x}_{t}|y)$
$=$ $-(\frac{1}{\alpha_{t}}\frac{\mathrm{d}\sigma_{t}}{\mathrm{d}t}-\frac{\sigma_{t}}{\alpha_{t}^{2}}\frac{\mathrm{d}\alpha_{t}}{\mathrm{d}t})\sigma_{t}\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x}_{t}|y)$
$=$ $\frac{\mathrm{d}(\sigma_{t}/\alpha_{t})}{\mathrm{d}t}\left(-\sigma_{t}\nabla_{\boldsymbol{x}}\log p_{t}(\boldsymbol{x}_{t}|y)\right)$
$=$ $\frac{\mathrm{d}(\sigma_{t}/\alpha_{t})}{\mathrm{d}t}\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t}|t,y).$
This equation is the scaled version of Diffusion PF-ODE under the notation of Karras et al. [karras2022elucidating].
0.D.2 DDIM Sampling
We derive the DDIM[song2020denoising] sampling algorithm in this section. According to the first order discretization of Eq. 6:
$\frac{\boldsymbol{x}_{t}}{\alpha_{t}}-\frac{\boldsymbol{x}_{s}}{\alpha_{s}}=(\frac{\sigma_{t}}{\alpha_{t}}-\frac{\sigma_{s}}{\alpha_{s}})\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{s}|s,y),$ (7)
the sampling algorithm of DDIM[song2020denoising] can be derived as the following equation:
$\boldsymbol{x}_{t}=\alpha_{t}(\frac{\boldsymbol{x}_{s}-\sigma_{s}\boldsymbol{\epsilon}_{\phi}(x_{s}|y,s)}{\alpha_{s}})+\sigma_{t}\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{s}|y,s).$ (8)
0.D.3 Proof of Proposition 1
We present a detailed derivation of Proposition 1 in the main text in this section. We define
$\boldsymbol{x}_{t}=\alpha_{t}\hat{\boldsymbol{x}}^{\text{c}}_{t}+\sigma_{t}\tilde{\boldsymbol{\epsilon}},$ (9)
[…]
$\frac{\mathrm{d}\hat{\boldsymbol{x}}^{\text{c}}_{t}}{\mathrm{d}t}=$ $\frac{\mathrm{d}\frac{\boldsymbol{x}_{t}-\sigma_{t}\tilde{\boldsymbol{\epsilon}}}{\alpha_{t}}}{\mathrm{d}t}$
$=$ $\frac{\mathrm{d}(\boldsymbol{x}_{t}/\alpha_{t})}{\mathrm{d}t}-\frac{\mathrm{d}(\sigma_{t}/\alpha_{t})}{\mathrm{d}t}\tilde{\boldsymbol{\epsilon}}$
$=$ $\frac{\mathrm{d}(\sigma_{t}/\alpha_{t})}{\mathrm{d}t}(\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t}|t,y)-\tilde{\boldsymbol{\epsilon}}).$
We also derive first-order discretization of FSD for 2D generation from Eq. 7:
$\hat{\boldsymbol{x}}^{\text{c}}_{t}-\hat{\boldsymbol{x}}_{s}^{o}=(\frac{\sigma_{t}}{\alpha_{t}}-\frac{\sigma_{s}}{\alpha_{s}})(\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{s}|s,y)-\tilde{\boldsymbol{\epsilon}}),$ (10)
Appendix 0.E Visualize the Change-of-Variable
(a) FSD (first-order discretization)
(b) DDIM[song2020denoising]
Figure 7: Generation results of FSD and DDIM. We apply FSD on 2D image generation using first-order discretization Eq. 10 instead of Adam[kingma2014adam]. In this case, we find FSD is the same as DDIM[song2020denoising] except some negligible differences, which may come from the differences on handling initial conditions.
Even though we show FSD can generate similar images When using the same initial noise in the main text, there are notable differences between the generated images since FSD uses Adam[kingma2014adam] to update parameters while DDIM uses first-order discretization of Diffusion PF-ODE. We show generation results of FSD that use first-order discretization in Fig. 7.
We also visualize the “change-of-variable” trick we used in the derivation of the main theorem. We define our new variable: the clean image $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ according to the following equation:
$\boldsymbol{x}_{t}=\alpha_{t}\hat{\boldsymbol{x}}^{\text{c}}_{t}+\sigma_{t}\tilde{\boldsymbol{\epsilon}}.$ (11)
Notably, there is also another similar but different concept: the one-step estimated ground-truth image $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$, defined by:
$\boldsymbol{x}_{t}=\alpha_{t}\hat{\boldsymbol{x}}^{\text{gt}}_{t}+\sigma_{t}\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t}|y,t).$ (12)
(a) noisy image $\boldsymbol{x}_{t}$
(b) clean image $\hat{\boldsymbol{x}}^{\text{c}}_{t}$
(c) one-step estimated ground-truth image $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$
Figure 8: Visualization of $\boldsymbol{x}_{t}$, $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ and $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$ in DDIM generation process. $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ is different from $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$. Moreover, one can see DDIM generation process as aligning $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ with $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$ since $\hat{\boldsymbol{x}}^{\text{gt}}_{t}-\hat{\boldsymbol{x}}^{\text{c}}_{t}\propto\nabla_{\theta}L_{\text{FSD}}^{\theta}$
(a) noisy image $\boldsymbol{x}_{t}$
(b) clean image $\hat{\boldsymbol{x}}^{\text{c}}_{t}$
(c) one-step estimated ground-truth image $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$
Figure 9: Visualization of $\boldsymbol{x}_{t}$, $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ and $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$ in FSD generation process. Compared to DDIM, FSD uses Adam[kingma2014adam] to update parameters. As a result, FSD needs more iterations to converge (See Tab. 3).
(a) noisy image $\boldsymbol{x}_{t}$
(b) clean image $\hat{\boldsymbol{x}}^{\text{c}}_{t}$
(c) one-step estimated ground-truth image $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$
Figure 10: Visualization of $\boldsymbol{x}_{t}$, $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ and $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$ in SDS generation process. Compared to FSD, SDS adds random noise instead of fixed noise. As a result, the estimated GT images vary a lot and SDS needs a larger learning rate to converge (See Tab. 3).
We visualize the trajectory of $\boldsymbol{x}_{t}$, $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ and $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$ in the same DDIM generation process in Fig. 8. As $\hat{\boldsymbol{x}}^{\text{gt}}_{t}-\hat{\boldsymbol{x}}^{\text{c}}_{t}=\frac{\sigma_{t}}{\alpha_{t}}(\tilde{\boldsymbol{\epsilon}}-\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t}|y,t))\propto\nabla_{\theta}L_{\text{FSD}}^{\theta}$ implies, we can see DDIM as a process that tries to align clean image with the one-step estimated ground-truth image generated by the Diffusion Model. We also visualize $\boldsymbol{x}_{t}$, $\hat{\boldsymbol{x}}^{\text{c}}_{t}$ and $\hat{\boldsymbol{x}}^{\text{gt}}_{t}$ of FSD and SDS in Fig. 9 and Fig. 10, respecitively.
Appendix 0.F Algorithm for Flow Score Distillation
We provide a summarized algorithm for Flow Score Distillation in Algorithm 1.
1:  Input: Text-to-image Diffusion Model $\boldsymbol{\epsilon}_{\phi}$ and prompt $y$. Learning rate $\eta$ for parameters of the 3D representation. A monotonically decreasing function $t(\tau)$. Blending factor $\beta$.
2:  Compute $\boldsymbol{\epsilon}_{b}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and $\boldsymbol{\epsilon}_{p}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$.
3:  for $\tau\in[0,\tau_{\text{end}}]$ do
4:     Randomly sample camera parameter $\boldsymbol{c}$.
5:     Render image $\boldsymbol{g}_{\theta}(\boldsymbol{c})$ and opacity mask $\boldsymbol{M}$ from 3D representation $\theta$.
6:     Randomly sample $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$.
7:     $\boldsymbol{\epsilon}(\boldsymbol{c})\leftarrow\sqrt{\beta}\cdot\left((1-\boldsymbol{M})\odot\boldsymbol{\epsilon}_{b}+\boldsymbol{M}\odot\boldsymbol{W}_{(W\frac{\phi_{\text{cam}}}{2\pi},H\frac{\theta_{\text{cam}}}{\pi})}(\boldsymbol{\epsilon}_{p})\right)+\sqrt{1-\beta}\cdot\boldsymbol{\epsilon}$.
8:     $\theta\leftarrow\theta-\eta\cdot(\boldsymbol{\epsilon}_{\phi}(\boldsymbol{x}_{t}|y,t)-\boldsymbol{\epsilon}(\boldsymbol{c}))\frac{\partial\boldsymbol{g}_{\theta}(\boldsymbol{c})}{\partial\theta}$
9:  end for
Algorithm 1 Flow Score Distillation
Appendix 0.G Practical Designing Rules for $\boldsymbol{\epsilon}(\boldsymbol{c})$
0.G.1 Practical Designing Rules
We do not specify the form of $\boldsymbol{\epsilon}(\boldsymbol{c})$ in the general form of FSD in the main text. However, it is intuitive to align $\boldsymbol{\epsilon}(\boldsymbol{c})$ in 3D space like the noise priors used in Video Diffusion Models[chang2024how, ge2023preserve, qiu2023freenoise]. We have tried several designs of $\boldsymbol{\epsilon}(\boldsymbol{c})$ and summarized several design rules for designing $\boldsymbol{\epsilon}(\boldsymbol{c})$ as well as the related potential problems if violating the design rules for $\boldsymbol{\epsilon}(\boldsymbol{c})$:
  • The noise associated with each camera view conditioning on camera view should be an uncorrelated Gaussian noise. i.e. $\boldsymbol{\epsilon}(\boldsymbol{c})|\boldsymbol{c}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. If not so, the input image may be out-of-distribution for the pretrained text-to-image Diffusion Models.
  • The noise associated with different camera views should be aligned in 3D space. Otherwise, FSD may degrade to the original SDS. This is consistent with the observation of recent work[ge2023preserve] on Video Generation, which showed that the noise maps corresponding to different video frames are highly correlated.
  • The noise should not be only aligned in specific points in the 3D space. This may lead to broken geometry since the convergence speed in 3D space is not uniform (see analysis on failure of constant noise function in the main text).
0.G.2 World-map Noise Function
We will take a deeper analysis of the property of the world map noise function in the main text. We will discuss how our proposed design rules are followed.
As noted in the main text, the local texture of the added noise is highly correlated with the generated images. In 3D spaces, it is natural to identify points, that correspond to similar points on the noise maps when projected onto the image planes of different camera views $\boldsymbol{c}$, as the points of high convergence speed. Let us denote those points as $\boldsymbol{p}_{+}$. For simplicity, we study the $\boldsymbol{p}_{+}$ that corresponds to the center points in image space.
For object-centric generation, a common camera view sampling strategy is to sample camera positions at $(r_{\text{cam}},\theta_{\text{cam}},\phi_{\text{cam}})$ under spherical coordinate and force the camera to look at the center point. When we apply a constant noise function as $\boldsymbol{\epsilon}(\boldsymbol{c})$, the center point in 3D space is a $\boldsymbol{p}_{+}$ point. As a result, FSD trends to generate geometry with holes. When applying world-map noise function proposed in main text, $\boldsymbol{p}_{+}$ is located at $(r_{+},\theta_{\text{cam}},\phi_{\text{cam}})$. One can compute that
$r^{\theta_{\text{cam}}}_{+}\approx\frac{2\tan\frac{\text{FOV}}{2}}{2\tan\frac{\text{FOV}}{2}+\Theta}\cdot r_{\text{cam}}$ (13)
if consider nearby views with slightly different $\theta_{\text{cam}}$ and
$r^{\phi_{\text{cam}}}_{+}\approx\frac{2\tan\frac{\text{FOV}}{2}}{2\tan\frac{\text{FOV}}{2}+\Theta\cdot\sin\theta_{\text{cam}}}\cdot r_{\text{cam}}$ (14)
if consider views with slightly different $\phi_{\text{cam}}$. In this way, we align nearby views coarsely. Moreover, for different camera parameters, $r_{+}$ is different, avoiding nonuniform convergence speed in 3D space.